You just opened a retirement account with a $5,000 deposit. Assuming that you can earn an 8% annual return (and that you make no additional deposits):

A. What will this account be worth when you retire in 35 years?

B. Now, assume that you wait 15 years before making the initial (and only) $5,000 deposit to your retirement account. What will your account be worth when you retire (note, you are still retiring 35 years from today)?

C. Why is the difference between your answers to A. and B. so large, and does this example suggest an optimal retirement strategy?

***To receive credit, you must show work!!!***
Answer #1

Answer A:

FV = PV * (1 + Periodic Interest rate) ^ Number of periods

Account will be worth when you retire in 35 years = FV = 5000 * (1 + 8%) ^35 = $73,926.72

Account will be worth when you retire in 35 years = $73,926.72

Answer B:

Assuming that you wait 15 years before making the initial (and only) $5,000 deposit, number of years amount remains invested = 35 - 15 = 20 years

Account will be worth when you retire in 35 years = 5000 * (1 + 8%) ^20 = $23304.79

Your account will be worth when you retire = $23,304.79

Answer C:

Difference between your answers to A and B is large since:

1. In scenario A, although same amounts is invested, it is invested early, when you have 35 years to retirement. As such the amount remained invested for 35 years.

2. The difference swells due to the compounding effect (interest earnings on interest amounts).

Optimal retirement strategy is to plan and start investing for retirement early. The impact of compounding affect is significant.